Cooperative San José of southern Sonora state in Mexico makes a unique syrup using cane sugar and local herbs. The syrup is sold in small bottles and is prized as a flavoring for drinks and for use in desserts. The bottles are sold for $12 each. The first stage in the production process is carried out in the Mixing Department, which removes foreign matter from the raw materials and mixes them in the proper proportions in large vats. The company uses the weighted-average method in its process costing system.

A hastily prepared report for the Mixing Department for April appears below:

Units to be accounted for:
Work in process, April 1 (materials 90% complete;
conversion 80% complete)
8,700
Started into production 37,600
Total units to be accounted for 46,300
Units accounted for as follows:
Transferred to next department 37,100
Work in process, April 30 (materials 75% complete;
conversion 50% complete)
9,200
Total units accounted for 46,300
Cost Reconciliation
Cost to be accounted for:
Work in process, April 1 $ 24,534
Cost added during the month 115,404
Total cost to be accounted for $ 139,938
Cost accounted for as follows:
Work in process, April 30 $ 19,734
Transferred to next department 120,204
Total cost accounted for $ 139,938

Management would like some additional information about Cooperative San José’s operations.

Required:

1. What were the Mixing Department's equivalent units of production for materials and conversion for April?

What were the Mixing Department's equivalent units of production for materials and conversion for April?

Materials Conversion
Equivalent units of production

2. What were the Mixing Department's cost per equivalent unit for materials and conversion for April? The beginning inventory consisted of the following costs: materials, $16,530; and conversion cost, $8,004. The costs added during the month consisted of: materials, $75,870; and conversion cost, $39,534.

Materials Conversion
Cost per equivalent unit

3. How many of the units transferred out of the Mixing Department in April were started and completed during that month?

Units started and completed during April

4. The manager of the Mixing Department stated, “Materials prices jumped from about $1.80 per unit in March to $2.30 per unit in April, but due to good cost control I was able to hold our mater

The manager of the Mixing Department stated, “Materials prices jumped from about $1.80 per unit in March to $2.30 per unit in April, but due to good cost control I was able to hold our materials cost to less than $2.30 per unit for the month.” Should this manager be rewarded for good cost control?

Yes or No
Answer #1

1. Equivalent units of Production = Units Transferred to next department + Ending Work in Process x Percentage Completion

Materials = 37100 + 9200 x 75% = 44000 units
Conversion = 37100 + 9200 x 50% = 41700 units

2.

Direct
Materials
Conversion
Costs
Total
a Beginning Work in Process $       16,530.00 $          8,004.00 $           24,534.00
b Costs added during month $       75,870.00 $        39,534.00 $        115,404.00
c Costs to be accounted for (a+b) $       92,400.00 $        47,538.00 $        139,938.00
d Total Equivalent Units $       44,000.00 $        41,700.00 $           85,700.00
e Cost per equivalent unit (c/d) $                  2.10 $                  1.14 $                     3.24

3. Units started and completed during April = Units transferred to next department - Beginning Work in Process
= 37100 - 8700 = 28400 units

4. No, the manager should not be rewarded for good cost control. The Mixing Department’s low unit cost for April occurred because the costs of the prior month have been averaged in with April’s costs.